OpenStudy (anonymous):

Find the solution to this system of equations.. -7x + 6y = 59 -8x + 6y = 58 @texaschic101

OpenStudy (anonymous):

@elephantine

OpenStudy (anonymous):

@thomaster @study-mode @eliassaab

OpenStudy (anonymous):

Do you have an idea of how to start?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Okay so you could use substitution or elimination. Which do you prefer?

OpenStudy (anonymous):

Wait don't you do this... -7x + 6y = 59 -8x + 6y = 58 -------------- add or something along them lines?

OpenStudy (anonymous):

Yes, but first you need to make one of the terms cancel out. So you could multiply -7x+6y=59 by -1. Which will result in 7x-6y=-59. Then when you add down the -6y and 6y will cancel out.